Please help me to solve these problems step by step. Thank You.
(3a+2b)cube
and
(3-a)cube

Answer by stanbon(75887)

Form: (x+y)^3 = x^3+3x^2y+3xy^2+y^3
------------------------------------------
(3a+2b)cube
= (3a)^3 + 3(3a)^2*(2b)+3(3a)*(2b)^2+(2b)^3
= 27a^3 + 54a^2b + 36ab^2+8b^3
---------------------------------------------
Form: (x-y)^3 = x^3-3x^2y+3xy^2-y^3
and
(3-a)cube = 3^3-3(3^2)(a)+3(3)(a^2)-a^3
= 27-27a+9a^2-a^3
====================
Cheers,
Stan H.